OpenStudy (anonymous):

Multiply and simplify if possible. sqrt 7x(sqrt x - 7 sqrt 7)

OpenStudy (anonymous):

\[\sqrt{7x}(\sqrt{x}-7\sqrt{7})\]

OpenStudy (anonymous):

hmm..

OpenStudy (cwrw238):

note:- sqrt7 * sqrt7 = 7 and sqrx * sqrtx = x

OpenStudy (anonymous):

I distributed it and got sqrt7x^2 - 7sqrt14x